# Leveraging Multi-source and Heterogeneous Signals for Fatigue Detection

### v2

@@ IV Experiments
A series of experiments are designed to validate the proposed target-enhancement framework. Section 4.1 details the experimental setup, including the selected dataset, environment, and data preprocessing, while Section 4.2 describes the experimental design and discusses the results.
@@ IV-A Experimental Setup
For the experiments to be elaborated on in a later subsection, three datasets containing various physiological signals-some overlapping and others unique-are selected. Meanwhile, three widely used classification methods are chosen, utilizing both individual datasets and the enhanced target data for training. The details of the selected datasets are in Subection 4.1.1. To minimize discrepancies and enhance the usability of these datasets, we perform data preprocessing as described in Subsection 4.1.2, following the introduction of the experimental environment.
@@ IV-A1 Datasets
Out of all publicly available datasets for fatigue, VPFD[40], MEFAR[52] and FatigueSet[62] are chosen for our experiments because they share a large set of overlapping physiological signals while each also contains unique ones. For instance, all of three datasets include HR, PPG, GSR, ST, and ACC. Additionally, MEFAR provides EEG data, FatigueSet includes both EEG and ECG data, and VPFD offers unique eye-tracking data.
  • VPFD dataset[40] is a multimodal fatigue detection dataset designed to reflect practical deployment conditions. It was collected using wearable devices like the HoloLens and Google Pixel Watch 2. The dataset contains physiological signals from four 27-year-old participants monitored for up to 9 hours during sports, driving, and research activities.
  • MEFAR dataset[52] focuses on occupational mental fatigue analysis. It comprises neurophysiological data from 23 participants across four occupational groups (i.e., academicians, technicians, engineers, and kitchen workers). The EEG signals were captured by the NeuroSky MindWave, while other signals were recorded by the Empatica E4 wristband.
  • FatigueSet dataset[62] contains data from 12 participants performing controlled physical and cognitive tasks. It utilizes the Muse S headband for EEG signals, the Zephyr BioHarness chestband for ECG signals, and the Empatica E4 wristband for the rest of physiological signals.
@@ IV-A2 Implementation Environment and Data Preprocessing
Experiments are executed on a high-performance PC featuring an Intel Core i9-14900K processor, NVIDIA RTX 4090 GPU, and 128GB of RAM. The system operates on Windows 11, with a software stack comprising Python 3.10.15, PyTorch 2.4.1, scikit-learn 1.1.3, pandas 2.2.2, and numpy 1.26.4.
To minimize noise and artifacts, we apply a series of tailored preprocessing techniques to each signal type. Specifically, ACC and PPG signals are processed using Recursive Least Squares (RLS) and Single-channel Signal Analysis (SSA) to effectively remove motion artifacts while preserving critical physiological features. HR, GSR, and ST signals undergo Maximum Outlier Filtering to eliminate extreme values and ensure data smoothness. Additionally, all signals are resampled to 32 Hz to standardize sampling rates across various devices.
For the FatigueSet dataset, a participant-specific normalization strategy is implemented. Each subject’s data is normalized individually to mitigate inter-subject physiological variability, ensuring that all target-domain samples reside in a unified feature space, thereby enhancing model comparability and training stability.
Before conducting the experiments, we preprocess the dataset and divide it into an 80% training set and a 20% test set using a block-based strategy rather than random splitting. The data is first segmented into continuous blocks based on changes in fatigue labels. For instance, if biometric data is collected from 15 participants, each participant’s data forms a block. From each block, we select 20% of the data for the test set by taking the first 10% and the last 10%. Combined, these selections constitute 20% of the entire dataset, preserving the temporal order and capturing all states without shuffling.
This approach preserves the temporal integrity of the time-series data while maintaining a balanced distribution of fatigue labels, thus ensuring that the test data remained independent of the training process. By preventing test data from being seen during training, this strategy allows for a fair and reliable assessment of model performance while also providing a robust foundation for subsequent model training.
@@ IV-B Experimental Design and Results
IV-B1 Regression‑Based Imputation Model Selection
Without loss of generality, the first set of experiments validates our imputer selection by assessing its ability to recover a single blocked modality. We compare three regression architectures—MLP, LSTM (Long Short-Term Memory), and CNN1D—on two datasets: FatigueSet (with ECG masked) and MEFAR (with EEG masked). In each case, the missing channel is replaced by the model prediction, and we record the mean squared error. Table II presents these results. The MLP consistently achieves the lowest loss, demonstrating its superiority as an imputer. Consequently, we employ the MLP for all subsequent modality imputation tasks.
TABLE II: Regression performance of different models
Model FatigueSet-ECG MEFAR-EEG
@@ IV-B2 Validity of Regression‑Imputed Data Versus Noise Baselines
The second set of experiments is designed to not only verify Theorems 1 and 2 but also demonstrate that imputed data can effectively support classifier tasks. In this case, VPFD is chosen with its gaze modality masked and reconstructed using three different methods: MLP-predicted values, MLP prediction with Gaussian noise, and pure random noise. We refer to the modified datasets as $\text{VPFD}_{\text{MLP}}$, $\text{VPFD}_{\text{MLP+noise}}$, and $\text{VPFD}_{\text{noise}}$, respectively. The added noise follows a zero‑mean Gaussian distribution whose standard deviation is twice the largest absolute value in the original VPFD gaze data. Finally, we train four separate MLP classifiers using 80% of the original VPFD dataset and each of the three modified versions, respectively. All models are evaluated on the same test set—20% of the original VPFD data—to ensure a consistent basis for comparison. Table III clearly shows that the classifier trained on regression-generated data maintains a similar performance to the one trained on the original data. This demonstrates the potential of data synthesis as a solution for addressing missing data. However, the introduction of noise, whether partial or complete, increases the distribution difference between the original and generated data, leading to degraded classifier performance.
TABLE III: Classification performance on different training datasets.
@@
IV-B3 Impact of Cross‑Domain Imputed Modalities on Classification Performance
The third experiments evaluate the VPFD dataset under four scenarios involving augmentation with reconstructed modalities derived from two auxiliary sources: FatigueSet and MEFAR. We begin by training the fatigue detector $f_{\phi}$ on the unmodified VPFD to establish baseline performance. Next, we run Algorithm 1 under three distinct configurations of auxiliary source availability: (a) only FatigueSet is available, which is used to impute the missing ECG modality, resulting in the variant referred to as VPFD w/ ECG; (b) only MEFAR is available, used to reconstruct the missing EEG channel, producing the variant termed as VPFD w/ EEG; and (c) both FatigueSet and MEFAR are available, enabling the sequential reconstruction of ECG and EEG for the VPFD w/ ECG+EEG variant. In all cases, the original VPFD features are preserved, and only the missing channels are imputed and appended. Again, each dataset is split, with 80% used to train a classifier and the remaining 20% used for evaluation. To assess the robustness and generalizability of our approach across model architectures, we train three different classifiers in this experiment—MLP, LSTM, and CNN1D—on each dataset variant. This setup allows us to evaluate not only how each additional reconstructed modality—individually or in combination—contributes to downstream fatigue detection performance, but also the effectiveness of our framework acorss diverse classifier types.
Table IV summarizes the classification accuracy of LSTM, CNN1D, and MLP models under the four scenarios. The results demonstrate that our targeted imputer effectively recovers missing modalities and that each added channel contributes enriches feature extraction, thereby boosting fatigue detection performance. Importantly, this feature‑level augmentation enables leveraging knowledge derived from high‑fidelity but deployment‑constrained sensors (e.g., ECG and EEG), thereby strengthening the performance of the practical sensor configuration (i.e., VPFD) without modifying its original features.
TABLE IV: Classification accuracy (%) under VPFD augmentation scenarios
Dataset MLP LSTM CNN1D
VPFD 81.47 83.61 60.34
VPFD w/ EEG 84.13 88.77 81.75
VPFD w/ ECG 84.14 86.99 81.55
VPFD w/ EEG+ECG 91.13 92.12 88.43
@@ IV-B4 Ablation of Batch Normalization and Jacobian Regularization
The fourth experiments, summarized in Table V, examines how different regularization strategies influence the quality of imputed modalities, ultimately, the resulting classification performance. The MLP classifier is used as the testbed in this experiment. An ablation study compares four settings—Batch Normalisation (BN) alone, Jacobian‑norm regularization alone, both combined, and neither (serving as the baseline). BN smooths the feature distribution, while the Jacobian term sharpens the network’s sensitivity to the subtle variations required for reconstructing missing channels. When applied together, these mechanisms complement each other: BN provides a stable representation space and the Jacobian constraint enhances gradient information, allowing the model to recover absent modalities most faithfully and to achieve the best overall accuracy. This complementary effect is particularly evident in the configuration where both auxiliary sources are available, as the inclusion of multiple reconstructed modalities (i.e., EEG+ECG) accentruates the performance gap bewteen models wih and without knowledge transfer. Although this study uses the MLP classifier for clarity of presentation, the underlying analysis indicates that the same synergy between distributional and gradient‑based constraints is expected to generalize to other classifiers.
TABLE V: Ablation study results (Accuracy %)
@@ V Conclusion
Fatigue detection is crucial in high-stake environments, where reduced alertness can have serious consequences. Despite advances in sensor technology, reliability and accessibility in real-world environments remain major challenges, as many high-end sensors, such as EEG, struggle with consistency, portability, and adaptability outside controlled settings. To address this, this paper proposes a proof-of-concept framework that enables knowledge transfer across differently instrumented sensor domains, aiming to enhance fatigue detection in real-world sensor-constrained scenarios. The experimental results provide empirical support for our hypothesis and validate the effectiveness and performance of this framework.
For proof-of-concept purposes, the current framework only employs simple regression-based imputers, which might not fully capture the intrinsic nonlinear relationships among modalities. Nonetheless, the current framework does not explicitly model interdependencies across multiple source domains; instead, it relies on a basic cascading process for multi-source enhancement. These limitations highlight important opportunities for future research, including the development of more expressive imputation techniques and principled multi-source modeling strategies to fully leverage cross-domain structure.

### v3

@@ IV Experiments
A series of experiments are designed to validate our framework. Section IV.A details the experimental setup, including the selected dataset, environment, and data preprocessing. Section IV.B describes the experimental design and discusses the results.
@@ IV-A Experimental Setup
For the experiments to be elaborated in a later subsection, three datasets containing various physiological signals some overlapping and others unique are selected. Meanwhile, three widely used classification methods are chosen, utilizing both individual datasets and the enhanced target data for training. To minimize discrepancies and enhance the usability of these datasets, data pre-processing is required before prior to experimentation. To that end, we describe the selected datasets, implementation environment, and the data pre-processing steps in detail.
@@ IV-A1 Datasets
Out of all publicly available datasets for fatigue, VPFD[10799435], MEFAR[DERDIYOK2024109896] and FatigueSet[kalanadhabhatta2021fatigueset] are chosen for our experiments because they share a large set of overlapping physiological signals while each also contains unique ones. For instance, all of three datasets include HR, PPG, GSR, ST, and ACC. Additionally, MEFAR provides EEG data, FatigueSet includes both EEG and ECG data, and VPFD offers unique eye-tracking data.
  • VPFD dataset[10799435] is a multimodal fatigue detection dataset designed to reflect practical deployment conditions. It was collected using wearable devices like the HoloLens and Google Pixel Watch 2. The dataset contains physiological signals from four 27 year old participants monitored for up to 9 hours during sports, driving, and research activities.
  • MEFAR dataset[DERDIYOK2024109896] focuses on occupational mental fatigue analysis. It comprises neurophysiological data from 23 participants across four occupational groups (i.e., academicians, technicians, engineers, and kitchen workers). The EEG signals were captured by the NeuroSky MindWave, while other signals were recorded by the Empatica E4 wristband.
  • FatigueSet dataset[kalanadhabhatta2021fatigueset] contains data from 12 participants performing controlled physical and cognitive tasks. It utilizes the Muse S headband for EEG signals, the Zephyr BioHarness chestband for ECG signals, and the Empatica E4 wristband for the rest of physiological signals.
@@ IV-A2 Implementation Environment and Data Preprocessing
Experiments are executed on a high performance PC featuring an Intel Core i9-14900K processor, NVIDIA RTX 4090 GPU, and 128GB of RAM. The system operates on Windows 11, with a software stack comprising Python 3.10.15, PyTorch 2.4.1, scikit-learn 1.1.3, pandas 2.2.2, and numpy 1.26.4.
To minimize noise and artifacts, we apply a series of tailored preprocessing techniques to each signal type. Specifically, ACC and PPG signals are processed using Recursive Least Squares (RLS) and Single channel Signal Analysis (SSA) to effectively remove motion artifacts while preserving critical physiological features. HR, GSR, and ST signals undergo Maximum Outlier Filtering to eliminate extreme values and ensure data smoothness. Additionally, all signals are resampled to 32 Hz to standardize sampling rates across various devices.
For all datasets, a participant-specific normalization strategy is implemented. Each subject’s data is normalized individually to mitigate inter-subject physiological variability, ensuring that all target-domain samples reside in a unified feature space, thereby enhancing model comparability and training stability.
Before conducting the experiments, we preprocess the dataset and divide it into an 80% training set and a 20% test set using a block-based strategy rather than random splitting. Specifically, the data is first segmented into continuous blocks based on changes in fatigue labels. For instance, if biometric data is collected from 15 participants, each participant’s data forms a block. From each block, we select 20% of the data for the test set by taking the first 10% and the last 10%. Combined, these selections constitute 20% of the entire dataset, preserving the temporal order and capturing all states without shuffling.
For all experiments described below, except for the sensitivity test, the learning parameters are initialized as listed in Table II.
TABLE II: Hyperparameters for Imputation and Classification
Parameter Imputation Phase Classification Phase
Learning rate 1e-3 3e-4
Batch size 512 256
Optimizer Adam Adam
Dropout rate - 0.2
Max epochs 2000 200
Early stopping patience 40 20
@@ IV-B Experimental Design and Results
IV-B1 Regression Based Imputation Model Selection
Without loss of generality, the first set of experiments validates our imputer selection by assessing its ability to recover a single blocked modality. We compare four regression architectures—MLP, LSTM (Long Short Term Memory), CNN1D, and TimeMixer[wang2024timemixerdecomposablemultiscalemixing] on two datasets: FatigueSet (with ECG masked) and MEFAR (with EEG masked). In each case, the missing channel is replaced by the model prediction, and we record the mean squared error (MSE). All errors are computed on strictly held-out, non-overlapping 10-s windows sampled at 32 Hz. The results illustrate that the MLP consistently yields the lowest MSE as shown in Table III. Consequently, the MLP serves as the default model for all subsequent modality imputation tasks, unless otherwise specified.
TABLE III: Regression performance of different models
Model FatigueSet-ECG(MSE) MEFAR-EEG(MSE)
@@ IV-B2 Validity of Regression‑Imputed Data Versus Noise Baselines
The second set of experiments is designed to not only verify Theorems 1 and 2 but also demonstrate that imputed data can effectively support classifier tasks. In this case, VPFD is chosen with its gaze modality masked and reconstructed using three different methods: MLP-predicted values, MLP prediction with Gaussian noise, and pure random noise. We refer to the modified datasets as $\text{VPFD}_{\text{MLP}}$, $\text{VPFD}_{\text{MLP+noise}}$, and $\text{VPFD}_{\text{noise}}$, respectively. The added noise follows a zero‑mean Gaussian distribution whose standard deviation is twice the largest absolute value in the original VPFD gaze data. Finally, we train four separate MLP classifiers using 80% of the original VPFD dataset and each of the three modified versions, respectively. All models are evaluated on the same test set—20% of the original VPFD data—to ensure a consistent basis for comparison. Table IV clearly shows that the classifier trained on regression-generated data maintains a similar performance to the one trained on the original data. This demonstrates the potential of data synthesis as a solution for addressing missing data. However, the introduction of noise, whether partial or complete, increases the distribution difference between the original and generated data, leading to degraded classifier performance.
TABLE IV: Classification performance on different training datasets.
@@
IV-B3 Impact of Cross-Domain Imputed Modalities on Classification Performance
The third set of experiments evaluates the VPFD dataset across different scenarios, where modalities are reconstructed from two auxiliary sources: FatigueSet and MEFAR using various augmentation methods and evaluated with multiple classifiers.
We first train the fatigue detector $f_{\phi}$ on the unmodified VPFD to obtain the baseline. Next, we run Algorithm 1 under three configurations of auxiliary source availability with the MLP and TimeMixer serving as imputers, respectively: (a) only FatigueSet is available, which is used to impute the missing ECG modality, resulting in the variant referred to as VPFD w/ ECG (MLP) and VPFD w/ ECG (TimeMixer); (b) only MEFAR is available, used to reconstruct the missing EEG channel, producing the variant termed as VPFD w/ EEG (MLP) and VPFD w/ EEG (TimeMixer); and (c) both FatigueSet and guanMEFAR are available, enabling the sequential reconstruction of ECG and EEG for the VPFD w/ ECG+EEG (MLP) and VPFD w/ ECG+EEG (TimeMixer) variants. In all cases, the original VPFD features are preserved, and only the missing channels are imputed and appended. Again, each dataset is split, with 80% used to train a classifier and the remaining 20% used for evaluation.
To assess the robustness and generalizability of our approach across model architectures, we train four different classifiers in this experiment—MLP, LSTM, CNN1D, and a Transformer—on each dataset variant. We assess performance using multiple metrics, including accuracy, binary F1, Area Under the ROC Curv(ROC AUC), Area Under the Precision–Recall Curve(PR AUC), macro F1, weighted F1, expected calibration error (ECE), and Brier score. This setup allows us to evaluate not only how each additional reconstructed modality—individually or in combination—contributes to downstream fatigue detection performance, but also the effectiveness of our framework across diverse classifier types.
Table V summarizes the classification metrics under these scenarios. The results demonstrate that our targeted imputers effectively recover missing modalities and that each added channel contributes to and enriches feature representations, thereby boosting fatigue detection performance. Importantly, this feature-level augmentation enables leveraging knowledge derived from high-fidelity but deployment-constrained sensors (e.g., ECG and EEG), thereby strengthening the performance of the practical sensor configuration (i.e., VPFD) without modifying its original features. It is interesting to note that when used individually, neither the TimeMixer, a more advanced imputer, nor the Transformer, a more advanced classifier, yields consistent improvements over conventional models. However, when applied together, they produce small but consistent gains across most configurations. This observation suggests that while traditional imputers and classifiers are sufficient to validate the effectiveness of our proof-of-concept framework, the combined use of advanced architectures provides a promising direction for future enhancement, which we will discuss further in the concluding section.
TABLE V: Classification metrics (%) under VPFD augmentation scenarios.
Dataset Metrics MLP LSTM CNN1D Transformer
[…]
@@ IV-B4 Ablation of Batch Normalization and Jacobian Regularization
The fourth experiments, summarized in Table VI, examines how different regularization strategies influence the quality of imputed modalities, ultimately, the resulting classification performance. The MLP classifier is used as the testbed in this experiment. An ablation study compares four settings—Batch Normalisation (BN) alone, Jacobian‑norm regularization alone, both combined, and neither (serving as the baseline). BN smooths the feature distribution, while the Jacobian term sharpens the network’s sensitivity to the subtle variations required for reconstructing missing channels. When applied together, these mechanisms complement each other: BN provides a stable representation space and the Jacobian constraint enhances gradient information, allowing the model to recover absent modalities most faithfully and to achieve the best overall accuracy. This complementary effect is particularly evident in the configuration where both auxiliary sources are available, as the inclusion of multiple reconstructed modalities (i.e., EEG+ECG) accentruates the performance gap bewteen models wih and without knowledge transfer. Although this study uses the MLP classifier for clarity of presentation, the underlying analysis indicates that the same synergy between distributional and gradient‑based constraints is expected to generalize to other classifiers.
TABLE VI: Ablation study results (Accuracy %)
@@
((a))
((b))
Figure 4: Sensitivity analysis of the Transformer on VPFD (EEG+ECG).
IV-B5 Sensitivity Tests
Building on the previous results, where the combination of TimeMixer as the imputer and Transformer as the classifier achieved the better gains, we perform sensitivity tests under this configuration. Specifically, we evaluate the Transformer on VPFD (EEG + ECG) augmented using both FatigueSet and MEFAR via TimeMixer. The goal is to assess model robustness to input perturbations and hyperparameter variations (Fig.4).
For noise robustness, we inject zero-mean Gaussian noise with varying intensity $\sigma$ into the standardized input features and fit a low-order regression of validation AUC against $\sigma$. The resulting curve remains nearly flat for small $\sigma$ values and decreases smoothly as $\sigma$ increases, indicating stability under mild perturbations and controlled degradation under stronger noise (Fig.4a).
For regularization, we vary the affine parameters $(\alpha,\beta)$ applied to standardized inputs as $x^{\prime}=\alpha x+\beta$, and fit a smooth response surface of AUC over $(\alpha,\beta)$. The surface exhibits a clear ridge at mid-range values, with performance declining only when $\alpha$ or $\beta$ deviates substantially from their nominal values (e.g., $\alpha>1$ or $|\beta|>0.2$), indicating robustness to moderate calibration differences but sensitivity to extreme scaling or shifting. The results indicate that the model remains stable under small scaling and bias variations, which are typical in real-world deployments where sensor calibration and preprocessing may differ.
@@ V Conclusion
Fatigue detection is crucial in high stakes environments, where reduced alertness can have serious consequences. Despite advances in sensor technology, reliability and accessibility in real world environments remain major challenges, as many high-end sensors, such as EEG, are expensive, highly sensitive to environmental conditions, and require substantial operational setup outside controlled settings. To address this challenge, this paper proposes a proof-of-concept framework that enables knowledge transfer across differently instrumented sensor domains, aiming to enhance fatigue detection in real-world sensor-constrained scenarios. In particular, our approach transfers knowledge learned from high-fidelity laboratory sensors to field-deployable systems by using shared-channel overlap as a bridge, thereby maintaining detection accuracy when some modalities are unavailable. Our experiments provide empirical support for our claim and validate the framework’s effectiveness and practicality, showing that transferred knowledge can substantially enhance detection when high-end sensors are unavailable.
The research can be extended in several directions. Although label heterogeneity is not present in our current case study, it remains an important factor for generalization in real-world applications. Differences in annotation subjectivity can introduce transfer bias and evaluation challenges. Thus approaches incorporating open-set recognition and label-shift calibration [10.1145/2768831] are worth exploring to mitigate such bias. In addition, our experiments show that using TimerMixer and Transformer yields small but consistent gains over other traditional configurations. We will further this exploration in our future research, developing jointly optimized imputation and classifier architectures to better capture nonlinear dependencies and enhance fatigue-detection performance. Finally, as the framework continues to evolve, we plan to examine its real-time deployment aspects, focusing on computational efficiency, latency, and resource requirements to ensure suitability for field applications.
[Uncaptioned image] LuoBin Cui (Member, IEEE)received the B.S. degree from Jiangsu University of Technology, P.R. China, in 2020, and the M.S. degree from Monmouth University, USA, in 2022. He is currently pursuing the Ph.D. degree in Electrical and Computer Engineering at Rowan University, USA. His research interests include machine learning, and artificial intelligence, with a particular focus on Graph Neural Networks and their applications. He has also conducted research on fatigue detection using physiological data and explored the development of educational games to enhance student comprehension in science, technology, and engineering fields.
[Uncaptioned image] Yanlai Wu received the B.S. degree in Artificial Intelligence from the School of Information Science and Engineering, Chongqing Jiaotong University, Chongqing, China. He is currently pursuing the M.S. degree in Electrical and Computer Engineering with the School of Engineering, Rowan University, USA. His current research interests include pattern recognition, computer vision, and natural language processing.
[Uncaptioned image] Ying Tang (Senior Member, IEEE) received the B.S. and M.S. degrees from the Northeastern University, P. R. China, in 1996 and 1998, respectively, and Ph. D degree from New Jersey Institute of Technology in 2001. She is currently Full Professor and the Undergraduate Program Chair of Electrical and Computer Engineering at Rowan University, Glassboro, New Jersey. Her current research interest lies in the area of cyber-physical social systems, extended reality, adaptive and personalized systems, modeling and adaptive control for computer-integrated systems, and sustainable production automation. Her work has been continuously supported by NSF, EPA, US Army, FAA, DOT, private foundations, and industry. She has three USA patents, and over 250 peer-reviewed publications, including 88 journal articles, 2 edited books, and 6 book/ encyclopedia chapters. Dr. Tang is presently Associate Editor of IEEE Transactions on Systems, Man, and Cybernetics: Systems, IEEE Transactions on Intelligent Vehicles, IEEE Transactions on Computational Social Systems, and Springer’s Discover Artificial Intelligence. She is the Founding Chair of Technical Committee on Intelligent Solutions to Human-aware Sustainability for IEEE Systems, Man, & Cybernetic, and the Founding Chair of Technical Committee on Sustainable Production Automation for IEEE Robotic and Automation.
[Uncaptioned image] Weikai Li received his B.S. degree in Information and Computing Science (2015) and M.S. degree in Computer Science and Technology (2018) from Chongqing Jiaotong University, Chongqing, China. He obtained his Ph.D. degree in Computer Science and Technology from the College of Computer Science and Technology, Nanjing University of Aeronautics and Astronautics, Nanjing, China, in 2022. He is currently a Lecturer at the School of Mathematics and Statistics, Chongqing Jiaotong University. He is also a Researcher with the MIIT Key Laboratory of Pattern Analysis and Machine Intelligence (PAMI), Nanjing, and China Science IntelliCloud Technology, Hefei, China. His research interests include pattern recognition and machine learning.